OpenStudy (anonymous):

64^(x-4)=4^2x can someone please explain to me how to solve this

OpenStudy (diyadiya):

\[\large 64= 4^3\]\[\Large ((4)^3)^{(x-4)}= 4^{2x}\]\[\Large 4^{3(x-4)}=4^{2x}\] Therefore 3(x-4) = 2x can you do it now?

OpenStudy (anonymous):

yeah but that first part confuses me can you tell me like what you did

OpenStudy (diyadiya):

ok 64 = 4^3 So inorder to make the bases same i replaced 64 with 4^3 Now we know that\[\Large (x^m)^n=x^{m \times n}\]Same way\[\Large (4^3)^{(x-4)} =4^{3(x-4)}\] Is this clear ?

OpenStudy (anonymous):

yeah thank you!

OpenStudy (diyadiya):

Anytime :)